40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Oct 30 01:56:38 2013, tara, DailyProgress, optic, table work oldPBS.jpgnewPBS.jpg
    Reply  Thu Oct 31 00:02:17 2013, tara, DailyProgress, Electronics Equipment, TTFSS offset.pngD040105C.png
       Reply  Thu Oct 31 18:10:07 2013, Evan, DailyProgress, Electronics Equipment, TTFSS 
       Reply  Sun Nov 3 19:37:32 2013, rana, DailyProgress, Electronics Equipment, TTFSS 
          Reply  Tue Nov 5 01:11:15 2013, tara, DailyProgress, Electronics Equipment, TTFSS 
             Reply  Wed Nov 6 01:14:58 2013, tara, DailyProgress, Electronics Equipment, TTFSS beat_2013_11_05.pngbeat_2013_11_05.fig
                Reply  Thu Nov 7 05:08:13 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_2013_11_07.pngbeat_2013_11_07.fig
                   Reply  Wed Nov 20 18:19:01 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_compare.pngbeat_compare.fig
Message ID: 1376     Entry time: Wed Oct 30 01:56:38 2013     Reply to this: 1377
Author: tara 
Type: DailyProgress 
Category: optic 
Subject: table work 

I'm optimizing the setup, and clearing the table a little bit.

  • Self homodyne setup in ACAV path is removed. This is from Erica's setup and it is not used. The input part is left, since I might use it for fiber distribution system
  • optics on RCAV path, all polarization are optimized. This includes, the input and output polarization for EOAM, and quarter wave plate before the periscope. The input polarization for sideband EOM is left intact after the last adjustment, and it should be good. With+/- 4V input, I can change the power by +/-10%, (1.0 +-0.1 mW is the current setup). For Evan: Do not touch anything before discussing with me!!!
  • I replaced a new PBS for PDH locking in RCAV path. The old one is bad. The surface between the prisms is milky, see the pictures below for comparison. There is also beams from multiple reflection within the cube. The new one is much better. There is no ghost beam anymore.
  • I blocked all the scattered light I could find in RCAV path with Irises and beam dumps. For ACAV, I just blocked the scattered lights from the laser to the PMC. I will finish the whole setup later.
  • I rechecked the height of the beam through EOMs/EOAMs. Since it is a little tricky to center the beam through the openings. The EOMs in RCAV path are all checked. For ACAV, only those between the laser and the PMC are checked(BB for phase locking and 21.5 for PMC sideband). The 14.75Mhz sideband and EOAM will be done later. The EOAM and wave plates are removed temporarily.
  • I modified the TTFSS for RCAV to have a gain reduction switch to help locking the laser. I tried to lock RCAV, but I cannot turn up the gain. I'm not sure what I did wrong but this has to be investigated.

To do lists

  • put optics back in ACAV path and optimize them (alignment + polarization).
  • fix RCAV TTFSS . Check by measuring the TF of the modified stage/ scanning laser + checking error signal

oldPBS.jpg

above: old PBS, bad inter surface can be seen.

newPBS.jpg

above: new PBS: all surfaces are clear

ELOG V3.1.3-